• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M.Shabestari

    Log entry time 00:04:51 on December22,2008

    Entry number 254345

    keyword=Swing shift summary

    Swing shift summary December 21, 2008

    SL : Mitra Shabestari
    TO : Abdurahim Rakhman

    =============================
    Target : 3He
    Beam current: 13 uA


    16:00 Shift started. Run 4664 has already started.
    Current HAPPEX run: 30614
    We continue taking production data.

    Run# | Start | Stop | Charge (u3)
    ------------------------------------------
    4664 | 15:41 | 16:40 | 0.04245
    4665 | 16:41 | 17:41 | 0.04214
    4666 | 17:43 | 18:44 | 0.04214
    4667 | 18:45 | 19:44 | 0.04251
    4668 | 19:45 | 20:44 | 0.01338

    20:15....Called MCC since beam has been down for ~12 minutes.
    He said they were having CAMAC crate problem, which is expected to be fixed
    in about 20 minutes.
    We stopped both runs. Yukai started some pedestal checks.

    20:55....Called MCC again.They are waiting for the expert to fix the problem. No beam.

    23:20....MCC called. They are ready to send beam. Moved target to empty. BPMA and BPMB look fine.Moved target to BeO.

    Run 4672 Spot++ run with BeO target to check the beam position.

    Moved target to 3He. Asked for 2uA and 5X6 raster.

    Noticed L14.5 in BigBite is off. Turned it back on. All previous runs look fine.
    There was a problem with DAQ.Rebooted CODA.

    Increased current to 5uA with 3X4 raster.

    Increased current to 13.5uA and started taking production data (Run 4676).

    Total charge collected during swing shift = 0.18262 C